OpenStudy (aliciarj):

You borrowed 250,000 from your parents to buy your first home. They are charging you 5% interest and the mortgage is for thirty years. Please prepare an Amortization schedule for the Mortgage for the first year of the loan.

OpenStudy (retireed):

It is easy to do with a calculator, but hard to do manually. I can give you the answer, but doesn't help you figure out how to do the problem. You need to figure out the totally monthly payment and then determine how much goes towards reducing the principal and how much goes towards paying the interest. How do they want you to solve the problem?

OpenStudy (aliciarj):

I just need help with starting it

OpenStudy (aliciarj):

for some reason i get stuck. Maybe if you can give me an example step by step

OpenStudy (wolf1728):

Here's the formula for monthly payment

OpenStudy (retireed):

Are you allowed to use a calculator or a website? Or are you suppose to use a mathematical formula to solve the problem? There are plenty of websites that can solve the this type of question.

OpenStudy (retireed):

@wolf1728 That is a good formula for solving the total payment, but how do you determine how much goes for principal reduction and how much goes towards interest?

OpenStudy (wolf1728):

If you go to this page http://www.1728.org/loanmath.htm it shows how to calculate interest, equity, etc. If you need a monthly calculator click here: http://www.1728.org/mortmnts.htm

OpenStudy (retireed):

To do it manually is pretty calculator intensive.

OpenStudy (wolf1728):

Monthly rate = .05/12 = 0.0041666667 Monthly Payment = ((0.0041666667 + 0.0041666667/[(1.0.0041666667)^360)-1]) * 250,000 =(0.0041666667 + 0.0012015496) * 250,000 = 1,342.054075 = 1.342.05 per month It's just that easy LOL

OpenStudy (wolf1728):

It seems as if aliciarj is offline.

OpenStudy (wolf1728):

Actually, it is pretty difficult to calculate the monthly payment (see above). Does your monthly amortization match my calculator? (It should).

OpenStudy (retireed):

I agree it is easy to calculate the TOTAL monthly payment, but it is difficult to calculate the amortization schedule for the first 12 months. It is not difficult, it is just time consuming to do it with a calculator. I just did the amortization schedule for the first two months in a spreadsheet and then it was easy to replicate for the first year. My spreadsheet and Quicken Loan payment matched up with my TI 84 total payment calculations,

OpenStudy (wolf1728):

Your balance after 1 year = 246,311.64 My total = $246,311.58 .06 difference (I guess mine works too)

OpenStudy (retireed):

Computer round off error. Close enough for government work.

OpenStudy (wolf1728):

Precisely!!

OpenStudy (aliciarj):

Thanks you guys

OpenStudy (aliciarj):

ok, I need to know how do you o from one month to the next payments are decreasing. PLease give me step by step for 2months. I want to know how to do it please

OpenStudy (wolf1728):

alicia I guess you came back

OpenStudy (aliciarj):

yes, was taking a test. I figured out how to do amortization schedule. Thanks